Hi, the decibel level of sound is given by D=10log(I/10^-12), where I is the sound intensity measured in watts per square meter. Find the decibel level of a whisper at an intensity of 5.4x10^-10 watts per square meter. can anyone help?
Answer by Nate(3500)

D = 10*log(I / 10^-12)
D = 10*log(5.4x10^-10 / 10^-12)
D = 10*log(5.4x10^-10x10^12)
D = 10*log(5.4x10^2)
D = 10*log(540)
About 27.3239